# Editable-DeepSC: Reliable Cross-Modal Semantic Communications for Facial Editing

### v3

@@ 3.1 Semantic Transmitter
The input image $S_{I}$ is first transformed by the Image Semantic Encoder into low-dimensional representations for data compression, i.e.,
@@
where $\mathcal{SE_{I}}(\cdot;\alpha)$ denotes the Image Semantic Encoder with the parameters $\alpha$. Similarly, the corresponding instructive sentence $S_{T}$ is also encoded by the Text Semantic Encoder $\mathcal{SE_{T}}(\cdot;\beta)$ with the parameters $\beta$ to acquire the textual semantic information, which is highly relevant to the anticipated editing:
@@
The output of $\mathcal{JECE}(\cdot,\cdot;\gamma)$, i.e., $X$, is then transmitted through the physical channels. We enforce an average transmission power constraint on $X$ to ensure that it does not exceed the limits of current communication capabilities:
@@
where $k$ is the length of $X$ and $P$ is the power constraint.
@@
$\rho=\frac{k}{H\times W\times C},$ (5)
where $H$, $W$, and $C$ represent the image’s height, width, and color channels. Smaller $\rho$ indicates better compression.
@@ 3.2 Semantic Receiver
where $Y$ represents the received edited semantic information, $h$ denotes the channel coefficients, and $N$ denotes the Gaussian noise, whose elements are independent of each other and have the same mean and variance.
Finally, $Y$ is mapped back to the high-dimensional vision domain to get the edited image required by the transmitter:
@@
where $\mathcal{JSCD}(\cdot;\theta)$ indicates the Joint Semantic-Channel Decoder with the parameters $\theta$. Note that $\mathcal{JSCD}(\cdot;\theta)$ will apply semantic decoding as well as channel decoding to $Y$ at the same time, which will reduce the semantic errors caused by the channel noise. Moreover, since only the edited images are ultimately required and there is no need to reconstruct the instructive sentences, the Text Semantic Decoder is unnecessary in our model and thus not considered.
@@ 2. $M_{2}$:
For the encoding and decoding steps of DNN as well as the channel transfer steps, the outputs of each stage are largely dependent on the current inputs instead of the previous outputs, which approximately satisfies the Markov property. Thus, as suggested in many previous works [38, 39, 40], we approximate the above encoding, decoding, and channel transfer steps as a Markov chain $K_{0}\rightarrow K_{1}\rightarrow\cdot\cdot\cdot\rightarrow K_{n}$, where $K_{0}$ indicates the distribution of original images at the transmitter side, $n$ is the total number of data processing procedures, $K_{n}$ indicates the distribution of ultimate edited images at the receiver side, and $K_{s}\;(0<s<n)$ indicates the distribution of a series of intermediate representations during the above steps. For any two transitions $U\rightarrow V\rightarrow W$ from the whole Markov chain, we have the Data Processing Inequality (DPI) [41] as demonstrated in Theorem 3.1.
Theorem 3.1.
@@
By applying the DPI to the whole Markov chain $K_{0}\rightarrow K_{1}\rightarrow\cdot\cdot\cdot\rightarrow K_{n}$ as pointed out in [38], we have $\mathcal{I}(K_{0};K_{1})\geq\mathcal{I}(K_{0};K_{2})\geq\cdot\cdot\cdot\geq\mathcal{I}(K_{0};K_{n})$. This implies that as the data processing procedures increase, the semantic mutual information about the original distribution $K_{0}$ carried by the intermediate distribution $K_{s}$ will decrease and cannot be recovered in the subsequent steps. Therefore, the aforementioned separate schemes $M_{1}$ and $M_{2}$ that have more encoding and decoding steps are more likely to lose semantic mutual information, which will eventually limit the perceptual performance of edited images. On the contrary, Editable-DeepSC integrates editings into the communication chain to decrease the data processing procedures, and will thus preserve more semantic mutual information with enhanced tasks execution performance.
@@ 4.1 Semantic Coding via Pre-trained GAN Inversion
We leverage the GAN inversion methods based on StyleGAN [25] priors to design the Image Semantic Encoder $\mathcal{SE_{I}}(\cdot;\alpha)$, as there is rich prior knowledge within the pre-trained GAN to reduce the representation dimension. To be specific, given an input image $S_{I}$, a random vector $z$ is first initialized in the StyleGAN latent space. The random vector $z$ is then fed into the pre-trained StyleGAN Generator $G(\cdot)$ to generate an initial image $I_{g}=G(z)$. The distortion $J(z)$ between $S_{I}$ and $I_{g}$ is calculated as:
@@
where $r_{inv}$ is the number of total inversion iterations, $z^{(0)}$ denotes the randomly initialized latent vector, $z^{(t)}\;(1\leq t\leq r_{inv})$ denotes the outcome after the $t$-th iteration, and $\eta_{inv}$ is the inversion learning rate. Upon completing the optimization of (9), we will obtain the output of semantic coding, which is denoted as $E_{I}=z^{(r_{inv})}$. Note that the dimension of $z$ is often much smaller than the dimension of $S_{I}$, hence we can greatly compress the original data and reduce the channel bandwidth ratio (CBR) defined in (5).
As for the Joint Semantic-Channel Decoder $\mathcal{JSCD}(\cdot;\theta)$, we first send $Y$ to another pre-trained StyleGAN Generator for semantic decoding to map the compressed data back to high-dimensional images. Besides, we additionally introduce lightweight trainable parameters for channel decoding to reduce the semantic errors caused by the channel noise, which we will detailedly discuss in Section 4.3.
4.2 Joint Editing-Channel Coding via Iterative Attributes Matching
@@
Joint Editing-Channel Coding (JECC). Inspired by the fact that DeepJSCC [26] combines source coding and channel coding to jointly optimize data compression and error correction, we implement the Joint Editing-Channel Encoder $\mathcal{JECE}(\cdot,\cdot;\gamma)$ by Fully Connected Layers $\mathcal{M}(\cdot;\gamma)$ to jointly optimize editings and communications for reduced data processing procedures. Note that $\mathcal{M}(\cdot;\gamma)$ only takes $E_{I}$ as the input, while $E_{T}$, the other input of $\mathcal{JECE}(\cdot,\cdot;\gamma)$, is utilized to obtain $i_{tar}$, $d_{tar}$ and $\Delta_{tar}$ that guide the semantic editing. In the test time, we first compute the expected degree of targeted attribute by $\{a_{1},a_{2},...,a_{i},...,a_{m}\}$, $i_{tar}$, $d_{tar}$, and $\Delta_{tar}$:
@@
where $r_{edit}$ is the maximum number of total editing iterations, $E_{I^{\prime}}^{(0)}$ is the same as $E_{I}$, and $E_{I^{\prime}}^{(t)}\;(1\leq t\leq r_{edit})$ denotes the outcome after the $t$-th iteration. After each iteration of modification, $E_{I^{\prime}}^{(t)}$ is delivered to the StyleGAN Generator $G(\cdot)$ to reconstruct an intermediate image, which will be sent to the pre-trained attribute predictor $P(\cdot)$ defined in [43] to check whether the anticipated requirement is satisfied. The predicted degree of the intermediate image calculated by $P(\cdot)$ will be compared with the expected degree $a_{exp}$. Once they match with each other, the iteration will stop and $E_{I^{\prime}}^{(t)}$ will be regarded as the output of $\mathcal{JECE}(\cdot,\cdot;\gamma)$. Otherwise, $\mathcal{M}(\cdot;\gamma)$ will continue to make small movements on $E_{I^{\prime}}^{(t)}$ until the target is reached or the maximum number $r_{edit}$ is exceeded.
@@
Figure 3: Illustration of our SNR-aware channel coding based on model fine-tuning. We introduce two lightweight trainable adapters that do not change the shapes of inputs and outputs to the Joint Editing-Channel Encoder and the Joint Semantic-Channel Decoder. When fine-tuning the models, only the parameters of these two adapters are adjusted to capture the distribution of new noise conditions, and the rest parameters are frozen to avoid forgetting the previously learned priors from Section 4.2.
@@ 4.3 SNR-aware Channel Coding via Model Fine-tuning
As shown in Figure 3, we introduce two lightweight trainable adapters that consist of Fully Connected Layers and do not change the shapes of inputs and outputs to the Joint Editing-Channel Encoder $\mathcal{JECE}(\cdot,\cdot;\gamma)$ and the Joint Semantic-Channel Decoder $\mathcal{JSCD}(\cdot;\theta)$. The output of $\mathcal{M}(\cdot;\gamma)$ is sent to the Channel Encoder Adapter $\mathcal{A}_{CE}(\cdot;\phi)$ to obtain the final outcome of channel encoding:
@@
where $r_{ft}$ is the number of total fine-tuning iterations, $\phi^{(t)},\psi^{(t)}\;(1\leq t\leq r_{ft})$ are the results after the $t$-th iteration, and $\eta_{ft}$ is the fine-tuning learning rate. When the fine-tuning is completed, these two adapters can capture the data distribution of current SNR level and enhance the perceptual quality of edited images in the test time, especially under extreme channel conditions. And the rest parameters are frozen to avoid forgetting the previously learned priors from Section 4.2.

### v4

@@ 3.1 Semantic Transmitter
The input image $S_{I}\in\mathbb{R}^{1\times C\times W\times H}$ (where $H$, $W$, and $C$ represent the image’s height, width, and color channels) is first transformed by the Image Semantic Encoder into low-dimensional representations for data compression, i.e.,
@@
where $E_{I}\in\mathbb{R}^{1\times L_{1}}$ (with $L_{1}$ being the dimension of visual features) is the extracted image semantic information and $\mathcal{SE_{I}}(\cdot;\alpha)$ denotes the Image Semantic Encoder with the parameters $\alpha$. Similarly, the corresponding instructive sentence $S_{T}\in\mathbb{R}^{1\times L_{2}\times L_{3}}$ (where $L_{2}$ is the text length and $L_{3}$ is the dimension of textual embeddings) is also encoded by the Text Semantic Encoder $\mathcal{SE_{T}}(\cdot;\beta)$ with the parameters $\beta$ to acquire the textual semantic information $E_{T}\in\mathbb{R}^{1\times L_{4}}$ (with $L_{4}$ being the dimension of textual features), which is highly relevant to the anticipated editing:
@@
The output of $\mathcal{JECE}(\cdot,\cdot;\gamma)$, i.e., $X\in\mathbb{C}^{1\times k}$ (where $k$ is the length of $X$), is then transmitted through the physical channels. We enforce an average transmission power constraint on $X$ to ensure that it does not exceed the limits of current communication capabilities:
@@
where $P_{max}$ is the power constraint.
@@
$\rho=\frac{k}{H\times W\times C}.$ (5)
Smaller $\rho$ indicates better compression. In our framework, the GAN inversion process (detailed in Section 4.1) plays a central role in achieving a low CBR. It compresses the high-dimensional input image $S_{I}\in\mathbb{R}^{1\times C\times W\times H}$ into a low-dimensional latent representation $E_{I}\in\mathbb{R}^{1\times L_{1}}$. This latent vector $E_{I}$ is further processed by the JECC encoder to constitute the signal $X\in\mathbb{C}^{1\times k}$ for transmission, whose length $k$ is also very small due to the previous compression step of GAN inversion, consequently resulting in smaller $\rho$.
@@ 3.2 Semantic Receiver
where $Y\in\mathbb{C}^{1\times k}$ represents the received signal, $h\in\mathbb{C}$ denotes the channel coefficients, and $N\in\mathbb{C}^{1\times k}$ denotes the Gaussian noise, whose elements (including the real and imaginary parts) are independent of each other and have the same mean and variance. Under simplified channel models such as the Additive White Gaussian Noise (AWGN) channel, $h$ is typically a scalar constant (e.g., $h=1$, representing no fading). In more general flat-fading channel models, $h$ is a complex random variable following a certain distribution (e.g., Rayleigh, Rician) that characterizes the attenuation and phase shift introduced by the wireless channel. For simplicity, we consider $h=1$ following the previous methods [55, 46].
Finally, $Y$ is mapped back to the high-dimensional vision domain to get the edited image $\hat{S}_{I^{\prime}}\in\mathbb{R}^{1\times C\times W\times H}$ required by the transmitter:
@@
where $\mathcal{JSCD}(\cdot;\theta)$ indicates the Joint Semantic-Channel Decoder with the parameters $\theta$. It first acts as a channel decoder, where its internal components (e.g., lightweight adapters) work to correct the errors introduced by the noisy channel, effectively denoising the received signal $Y$. Subsequently, it functions as a semantic decoder and maps the partially corrected latent representations back to the high-dimensional image space, reconstructing the final edited image $\hat{S}_{I^{\prime}}$. This integrated process ensures that semantic recovery and channel error correction are jointly optimized, rather than being treated separately, which will reduce the semantic errors caused by the channel noise. Moreover, since only the edited images are ultimately required and there is no need to reconstruct the instructive sentences, the Text Semantic Decoder is unnecessary in our model and thus not considered.
@@ 2. $M_{2}$:
For the encoding and decoding steps of DNN as well as the channel transfer steps, the outputs of each stage are largely dependent on the current inputs instead of the previous outputs, which approximately satisfies the Markov property. Thus, as suggested in many previous works [41, 36, 37], we approximate the above encoding, decoding, and channel transfer steps as a Markov chain $K_{0}\rightarrow K_{1}\rightarrow\cdot\cdot\cdot\rightarrow K_{n}$, where $K_{0}$ indicates the distribution of original images at the transmitter side, $n$ is the total number of data processing procedures, $K_{n}$ indicates the distribution of ultimate edited images at the receiver side, and $K_{s}\;(0<s<n)$ indicates the distribution of a series of intermediate representations during the above steps. For any two transitions $U\rightarrow V\rightarrow W$ from the whole Markov chain, we have the Data Processing Inequality (DPI) [9] as demonstrated in Theorem 1.
Theorem 1.
@@
By applying the DPI to the whole Markov chain $K_{0}\rightarrow K_{1}\rightarrow\cdot\cdot\cdot\rightarrow K_{n}$ as pointed out in [41], we have $\mathcal{I}(K_{0};K_{1})\geq\mathcal{I}(K_{0};K_{2})\geq\cdot\cdot\cdot\geq\mathcal{I}(K_{0};K_{n})$. This implies that as the data processing procedures increase, the semantic mutual information about the original distribution $K_{0}$ carried by the intermediate distribution $K_{s}$ will decrease and cannot be recovered in the subsequent steps. Therefore, the aforementioned separate schemes $M_{1}$ and $M_{2}$ that have more encoding and decoding steps are more likely to lose semantic mutual information, which will eventually limit the quality of edited images. On the contrary, Editable-DeepSC integrates editings into the communication chain to decrease the data processing procedures, and will thus preserve more semantic mutual information with enhanced tasks performance.
Figure 2: Overview of the proposed framework. Editable-DeepSC mainly consists of the Text Semantic Encoder, the Image Semantic Encoder, the Joint Editing-Channel Encoder, and the Joint Semantic-Channel Decoder, where channel noise corruptions from the real world are also considered.
@@ 4 Method
This section elaborates the proposed Editable-DeepSC framework, whose overall pipeline and corresponding three-stage data flow are illustrated in Figure 3. The framework operates through three cohesive stages: (1) Semantic Coding via Pre-trained GAN Inversion; (2) Joint Editing-Channel Coding (JECC) via Iterative Attributes Matching; (3) SNR-aware Channel Coding via Model Fine-tuning.
In the first stage, the input facial image $S_{I}$ and textual instruction $S_{T}$ are processed by the Image Semantic Encoder $\mathcal{SE}_{I}(\cdot;\alpha)$ and Text Semantic Encoder $\mathcal{SE}_{T}(\cdot;\beta)$, to extract compact latent representations $E_{I}$ and $E_{T}$ via GAN inversion. These representations then flow into the second stage, where the core JECC mechanism iteratively refines $E_{I}$ guided by $E_{T}$ to produce the transmitted signal $X$, simultaneously achieving semantic editing and channel coding. After transmission through a noisy channel $Y=h\ast X+N$, the third stage employs SNR-aware fine-tuning, where lightweight adapters within $\mathcal{JECE}(\cdot,\cdot;\gamma)$ and $\mathcal{JSCD}(\cdot;\theta)$ are selectively updated to adapt to varying channel conditions, culminating in the generation of the final edited image $\hat{S}_{I^{\prime}}$ by $\mathcal{JSCD}(\cdot;\theta)$. This integrated pipeline ensures precise execution of editings while maintaining robustness against channel noise, significantly reducing redundant data processing procedures compared to separate approaches.
@@ 4.1 Semantic Coding via Pre-trained GAN Inversion
We leverage the Generative Adversarial Networks (GAN) inversion methods [15, 51, 57, 13, 12, 32, 33] based on StyleGAN [27] priors to design the Image Semantic Encoder $\mathcal{SE_{I}}(\cdot;\alpha)$, as there is abundant prior knowledge within the pre-trained GAN to reduce the representation dimension. To be specific, given an input image $S_{I}$, a random vector $z$ is first initialized in the StyleGAN latent space. The random vector $z$ is then fed into the pre-trained StyleGAN Generator $G(\cdot)$ to generate an initial image $I_{g}=G(z)$. The distortion $J(z)$ between $S_{I}$ and $I_{g}$ is calculated as:
@@
where $r_{inv}$ is the number of total inversion iterations, $z^{(0)}$ denotes the randomly initialized latent vector, $z^{(t)}\;(1\leq t\leq r_{inv})$ denotes the outcome after the $t$-th iteration, and $\eta_{inv}$ is the inversion learning rate. The learning rate $\eta_{inv}$ is set to $0.10$ to ensure stable convergence, and the maximum number of iterations $r_{inv}$ is set to $110$ based on a trade-off between reconstruction quality and computational cost, as analyzed in Section 5.5. The optimization process terminates when either the maximum iteration count $r_{inv}$ is reached or the relative change in the distortion $J(z)$ between consecutive iterations falls below a threshold $\epsilon=1\times 10^{-5}$, i.e., $|J(z^{(t)})-J(z^{(t-1)})|/|J(z^{(t-1)})|<\epsilon$. Upon completing the optimization of (9), we will obtain the output of semantic coding, which is denoted as $E_{I}=z^{(r_{inv})}$. Note that the dimension of $z$ is often much smaller than the dimension of $S_{I}$, hence we can greatly compress the original data and reduce the Channel Bandwidth Ratio (CBR) defined in (5).
As for the Joint Semantic-Channel Decoder $\mathcal{JSCD}(\cdot;\theta)$, we first convert $Y$ back to a real-valued latent representation, which is then sent to another pre-trained StyleGAN Generator for semantic decoding to map the compressed data back to high-dimensional images. Besides, we additionally introduce lightweight trainable parameters for channel decoding to reduce the semantic errors caused by the channel noise, which we will detailedly discuss in Section 4.3.
Figure 3: Illustration of our SNR-aware channel coding based on model fine-tuning. We introduce two lightweight trainable adapters that do not change the shapes of inputs and outputs to the Joint Editing-Channel Encoder and the Joint Semantic-Channel Decoder. When fine-tuning the models, only the parameters of these two adapters are adjusted to capture the distribution of new noise conditions, and the rest of the parameters are frozen to avoid forgetting the previously learned priors from Section 4.2.
4.2 Joint Editing-Channel Coding (JECC) via Iterative Attributes Matching
@@
Joint Editing-Channel Coding (JECC). Inspired by the fact that DeepJSCC [5] combines source coding and channel coding to jointly optimize data compression and error correction, we implement the Joint Editing-Channel Encoder $\mathcal{JECE}(\cdot,\cdot;\gamma)$ by Fully Connected Layers $\mathcal{M}(\cdot;\gamma)$ to jointly optimize editings and communications for reduced data processing procedures. The Joint Editing-Channel Encoder $\mathcal{JECE}(\cdot,\cdot;\gamma)$ will iteratively exert minor modifications on the previous results of semantic coding to realize semantic editing and channel coding at the same time. Note that $\mathcal{M}(\cdot;\gamma)$ only takes $E_{I}$ as the input, while $E_{T}$, the other input of $\mathcal{JECE}(\cdot,\cdot;\gamma)$, is utilized to obtain $i_{tar}$, $d_{tar}$ and $\Delta_{tar}$ that guide the semantic editing. In the test time, we first compute the expected degree of targeted attribute by $\{a_{1},a_{2},...,a_{i},...,a_{m}\}$, $i_{tar}$, $d_{tar}$, and $\Delta_{tar}$:
@@
where $r_{edit}$ is the maximum number of total editing iterations, $E_{I^{\prime}}^{(0)}$ is the same as $E_{I}$, and $E_{I^{\prime}}^{(t)}\;(1\leq t\leq r_{edit})$ denotes the outcome after the $t$-th iteration. After each iteration of modification, $E_{I^{\prime}}^{(t)}$ is delivered to the StyleGAN Generator $G(\cdot)$ to reconstruct an intermediate image, which will be sent to the pre-trained attribute predictor $P(\cdot)$ defined in [23] to check whether the anticipated requirement is satisfied. The predicted degree of the intermediate image calculated by $P(\cdot)$ will be compared with the expected degree $a_{exp}$. Once they match with each other, the iteration will stop and $E_{I^{\prime}}^{(t)}$ will be used to form the complex-valued signal $X$ (e.g., by treating the first and second halves as the real and imaginary parts) for transmission. Otherwise, $\mathcal{M}(\cdot;\gamma)$ will continue to make small movements on $E_{I^{\prime}}^{(t)}$ until the target is reached or the maximum number $r_{edit}$ is exceeded.
@@ 4.3 SNR-aware Channel Coding via Model Fine-tuning
Benefiting from the recent successful applications of transfer learning techniques [7, 6], we propose SNR-aware channel coding based on model fine-tuning. As shown in Figure 3, we introduce two lightweight trainable adapters that consist of Fully Connected Layers and do not change the shapes of inputs and outputs to the Joint Editing-Channel Encoder $\mathcal{JECE}(\cdot,\cdot;\gamma)$ and the Joint Semantic-Channel Decoder $\mathcal{JSCD}(\cdot;\theta)$. The complex-valued signal $X$ is sent to the Channel Encoder Adapter $\mathcal{A}_{CE}(\cdot;\phi)$ to obtain the final outcome of channel encoding:
@@
where $r_{ft}$ is the number of total fine-tuning iterations, $\phi^{(t)},\psi^{(t)}\;(1\leq t\leq r_{ft})$ are the results after the $t$-th iteration, and $\eta_{ft}$ is the fine-tuning learning rate. When the fine-tuning is completed, these two adapters can capture the data distribution of current SNR level and enhance the perceptual quality of edited images in the test time, especially under extreme channel conditions. And the rest of the parameters are frozen to avoid forgetting the previously learned priors from Section 4.2 because of the fine-tuning. The algorithm framework of Editable-DeepSC is summarized in Algorithm 1.
Algorithm 1 Editable-DeepSC Training Pipeline
0: Training dataset $\mathcal{D}$; Pre-trained models $G(\cdot)$, $P(\cdot)$, $F(\cdot)$; Hyper-parameters.
0: Trained parameters $\gamma,\theta,\phi,\psi$.
1:Stage 1: Semantic Coding via GAN Inversion
2:Goal: Compress image $S_{I}$ to latent code $E_{I}$.
3: Encode each $S_{I}\in\mathcal{D}$ by optimizing $z$ to minimize distortion $J(z)$ between $S_{I}$ and $G(z)$ as in (9).
4: Obtain semantic code $E_{I}=z^{(r_{inv})}$.
5:Stage 2: JECC Training via Iterative Attributes Matching
6:Goal: Learn joint editing-channel coding.
7: Encode text $S_{T}$ to get editing targets $(i_{tar},d_{tar},\Delta_{tar})$ via $\mathcal{SE_{T}}(\cdot;\beta)$.
8: Iteratively refine $E_{I}$ using $\mathcal{M}(\cdot;\gamma)$ guided by editing targets as in (11).
9: Train $\gamma,\theta$ by minimizing $\mathcal{L}=\lambda_{pred}\mathcal{L}_{pred}+\lambda_{id}\mathcal{L}_{id}+\lambda_{disc}\mathcal{L}_{disc}$ under high SNRs.
10:Stage 3: SNR-aware Model Fine-tuning
11:Goal: Adapt to varying channel conditions.
12: For each target SNR, fine-tune adapters $\mathcal{A}_{CE}(\cdot;\phi)$ and $\mathcal{A}_{CD}(\cdot;\psi)$ as in (18).
13: Update $\phi,\psi$ by minimizing LPIPS loss $\mathcal{L}_{ft}$ on a small image set. Freeze the other parameters.